# The Indian Pulsar Timing Array Data Release 2: II. Customised Single-Pulsar Noise Analysis and Noise Budget

### v1

@@ 1 Introduction
Millisecond pulsars (MSPs) are rapidly spinning neutron stars with extraordinary rotational stability and act as highly accurate cosmic clocks (LorimerKramer2004; Hobbs+2019). Pulsar timing is a powerful technique that involves careful monitoring of the precise rotational behavior of MSPs by comparing the measured pulse Times of Arrival (ToAs) with timing models that encapsulate the pulsar’s rotational, astrometric and orbital properties (Edwards+2006). Gravitational Waves (GWs) passing across the line-of-sight to a pulsar modulate the ToAs of the pulsar measured at a terrestrial observatory (Estabrook+1975). The ultra-low frequency (nHz) GWs can be produced by populations of inspiraling Supermassive Black Hole Binaries (SMBHBs), cosmic string networks (Kibble1976), relic post-inflation GWs (Lasky2016), cosmological phase transitions (Grishchuk2005) or exotic dark matter species (Burke-Spolaor+2019). The incoherent superposition of these nHz GWs in the universe produces a stochastic Gravitational Wave Background (GWB: Manchester2006; Mingrelli+2017; Burke-Spolaor+2019). This extremely feeble background introduces tiny delays in the ToAs that depend on the underlying population properties of its progenitor. For an isotropic distribution of a population of circular inspiraling SMBHBs, the GWB has a characteristic spectral signature and manifests itself as quadrupolar spatial correlations described by the Hellings-Downs (HD) curve (Hellings1983), and depends only on the relative sky separations of pulsar pairs in the pulsar ensemble used for this purpose.
@@
Due to the high rotational stability of MSPs, such signatures can be detected by precise timing of these pulsars through Pulsar Timing Arrays (PTAs: Sazhin1978; Foster1990; Taylor2015). PTAs monitor an ensemble of MSPs and leverage the spatially correlated modulations in ToAs to detect and possibly characterise the GWB. However, this signature can be obscured by a variety of noise sources which may be intrinsic to the pulsar, or could be picked up as the pulsed radio emission propagates through the ionized interstellar medium (IISM) before it is received by the terrestrial observatories (Shannon2010; Donner2020). Careful modeling and characterisation of these noise sources, therefore, becomes central to the detection of a GWB. This pursuit of unveiling the existence of a GWB has led to the formation of PTA consortia across the globe, such as the European Pulsar Timing Array (EPTA: Stappers2006; Janssen2008), the Parkes Pulsar Timing Array (PPTA: Hobbs2013; Manchester2006), the North American Nanohertz Observatory for Gravitational Waves (NANOGrav: McLaughlin2013), the Indian Pulsar Timing Array (InPTA: Joshi+2018; Joshi2022), the MeerKAT Pulsar Timing Array (MPTA: Miles2023), and the Chinese Pulsar Timing Array (CPTA: 2016ASPC50219L). Collectively, these PTAs (except the CPTA) form the International Pulsar Timing Array (IPTA: Verbiest2016; Perera2019) consortium, which combines data and resources of the constituent PTAs to enhance the sensitivity to detect these nHz GWs.
@@
In an earlier study, Chen2021; Arzoumanian2020; Goncharov2021; Antoniadis2022 reported the presence of a Common Red Noise (CRN) process in multiple PTA datasets. However, this signal could not be characterised as a GWB due to the absence of statistically significant spatial correlations, even though the process showed temporal correlations, necessary for a common red process. More recently, the NANOGrav, PPTA and EPTA+InPTA analyses indicated the presence of a spatially correlated GWB signal (NG15yrGWB; PPTADR3GWB; EPTA+InPTA-III), apart from a high significance detection of CRN, bringing the field of ultra-low frequency GW astronomy into the limelight. However, accurate noise modeling is critical to bring out these spatial correlations and estimate the parameters of the spectral signature of the GWB.
@@
The present study focuses on the customised Single-Pulsar Noise Analysis (SPNA) of 27 MSPs from the InPTA Data Release 2 (Rana2025), building up on the work presented in (Srivastava2023), employing the methods of Bayesian inference to detect and quantify the various noise processes underlying PTA datasets. The InPTA dataset has the niche characteristic of simultaneous dual-band observations, which gives it a unique standpoint amongst global PTAs for precise modeling of chromatic processes. The rest of the paper is structured as follows. In Section 2, we give a brief description of the InPTA-DR2 observing strategy, timing analysis and data products. In Section 3, we discuss the noise processes and their modeling in detail, where in Section 3.2 we give the theoretical background of noise models and in Section 3.3 we discuss the analysis methodology adopted in the present work. We present the results in Sections 4 and 5, followed by a discussion in Section 6. We conclude with a brief summary in Section 7 and outline the future prospects related to this work in Section 8.
@@ 2 Dataset
The InPTA-DR2 dataset consists of 27 MSPs observed with a cadence of $\sim$$15$ days using the upgraded Giant Metrewave Radio Telescope (uGMRT: Swarup1991; gak+2017), which enables simultaneous multi-band observations. The dataset spans $\sim$$7.2$ years of total time baseline ($T_{\text{span}}$) – nearly twice that of InPTA DR1 (Tarafdar+2022), and includes concurrent observations in the $300-500\,\text{MHz}$ (Band 3) and $1260-1460\,\text{MHz}$ (Band 5) frequency ranges, with $100\,\text{MHz}$ or $200\,\text{MHz}$ bandwidths in each band, depending on the observing epoch. This leads to a maximum of 7 different backend-receiver configurations for the longest time baseline pulsars. The raw data was recorded with the GMRT Wideband Backend and observations employed the phased-array (PA) beam and the real-time coherent de-dispersion pipeline (CDP) in one or both bands, depending on the observing cycle (see Table 1 of Rana+2025). The ‘Pipeline for the Indian Pulsar Timing Array’ (PINTA: smj+2021) is used to perform RFI mitigation and partial folding, to convert the uGMRT raw data into a standard pulsar archive format (PSRFITS: hvm+2004) with full frequency resolution and several sub-integrations of $10\,\text{s}$ duration. Further technical details of InPTA observations and data reduction are provided in Sections 2 & 3 and illustrated in Figures 1 & 2 of Rana+2025.
@@
To generate noise-free templates in each frequency band – (i) high signal-to-noise ratio (S/N) observations (free of artifacts) are selected in both bands, (ii) a preliminary alignment of the pulse profiles is performed using an initial fiducial dispersion measure [^1] The integrated free electron column density along the line-of-sight to a pulsar is called the Dispersion Measure (DM), given by : $\text{DM}=\int_{\ell}n_{e}\,\mathrm{d}l$ (DM), (iii) the bandpass shape is corrected by equalising the off-pulse RMS across frequency channels, and (iv) optimal wavelet smoothing is applied using the psrsmooth tool of PSRCHIVE (hvm+2004). The final templates are then de-dispersed using a high-precision fiducial DM obtained through an iterative procedure. This approach minimises pulse-profile smearing in template profiles due to uncorrected DM delays. Furthermore, a refined procedure is employed to determine the optimal number of frequency sub-bands to partially collapse the data across frequency for each pulsar in the two bands. This optimisation balances two requirements – (i) to provide sufficient S/N and similar ToA precision in each sub-band, and (ii) to minimise radio-frequency dependent profile evolution. Details about these aspects are provided in Sections 4.1–4.3 of Rana+2025.
@@
Given the noise-free templates, optimised number of sub-bands, and fiducial DM estimates for each pulsar, the DMCalc (kmj+2021) pipeline is used to estimate frequency-resolved ToAs and DMs, along with DMXs [^2] DMXs are defined as the differences between the estimated DMs and fiducial DM, and correspond to a linear piece-wise fit to the DM model in time-domain, to account for temporal variations in the DM. computed from the estimated and fiducial DMs for every epoch. In the deterministic timing analysis, these epoch-wise DMXs are used to model dispersion delays. ToAs with large uncertainties or having poor S/N in a given sub-band and/or affected by radio-frequency interference (RFI) are flagged. Finally, the timing model parameters are fitted using the TEMPO2 package (Tempo2). Following these procedures, the resulting InPTA-DR2 dataset spans MJD 577686039957768-60399 (uGMRT observation cycles 314531-45, excluding cycle $36$), where the exact temporal coverage differs for each pulsar. Further details of the DMCalc pipeline and timing analysis are provided in Sections 4.4 and 5 of Rana+2025, respectively. The final InPTA-DR2 data products included Band $3+5$ and Band $3$ DM time series, par files containing the deterministic timing solutions with DMXs, and tim files containing the cleaned ToAs [^3] https://github.com/inpta/InPTA.DR2.git.
@@ 3 Noise modeling methodology
ToAs obtained from the pulsed emission of pulsars depend on a variety of deterministic and stochastic processes. The deterministic contributions are derived from the rotational, astrometric and orbital dynamics of the pulsar, the orbital motion of the Earth, as well as the propagation effects of the interstellar medium. These deterministic terms are captured by the timing model of the pulsar derived from the pulsar timing ephemeris. However, the stochastic components cannot be accurately accounted for, based on this deterministic timing model alone. These components include achromatic (such as spin noise) as well as chromatic (such as DM noise) processes that add to the overall measurement noise in the dataset. Inaccurate modeling of any of these noise processes may lead to false detections and incorrect characterisation of potential GW signals. Therefore, precise modeling of these processes is of immense importance in PTA data analyses to isolate the astrophysical signals of interest (DiMarco+2025). In this section, we describe, in detail, the theoretical construct of various deterministic and stochastic noise processes and the customised SPNA methodology employed to estimate their contributions to the overall InPTA-DR2 noise budget.
@@ 3.1 The Deterministic Timing Model
The deterministic timing model essentially provides a mapping of a ToA measured at the terrestrial observatory to a fiducial point on the emitted pulse from a pulsar by performing a series of transformations from the pulsar frame to an inertial frame at the Solar System Barycenter (SSB). These transformations include clock corrections to convert the local observatory time to the Universal Time ($\Delta_{\text{clock}}$: UTC), account for the arrival time delays due to differences in the topocentric and SSB reference frames ($\Delta_{\odot}$: DE440), the constant light-travel time between the pulsar and the Earth ($\Delta_{\gamma}$), propagation through the IISM ($\Delta_{\text{IISM}}$), the astrometric and binary orbital motion of the pulsar ($\Delta_{\text{binary}}$). Therefore, the observed arrival time ($t_{\text{obs}}$) of the pulse can be related to the emission time ($t_{\text{em}}$) as:
@@
These delays are incorporated in the deterministic timing model in the form of a par file which is then used to perform standard timing analysis using packages such as TEMPO2 (Tempo2) and PINT (Pint) to generate a best-fit timing solution. This solution is then used to derive the ToA (or timing) residuals which are essentially the differences in the observed ($t_{\text{obs}}$) ToAs and the predicted ToAs as per the best-fit timing solution ($\bar{t}_{\text{obs}}$). The additive terms $\mathcal{N}$ in equation 1 correspond to the residual noise in the measured ToAs, which can arise due to noise in the radiometer or intrinsic fluctuation of the pulsed emission ($\mathcal{N}_{\text{r}}$) or other stochastic noise processes ($\mathcal{N}_{\text{s}}$), such as the spin noise or DM noise, etc., that are not modeled completely by the deterministic timing model of the pulsar.
@@ 3.2 The Noise Processes
Based on the temporal properties of noise, the single-pulsar noise processes are categorised as – (i) white noise processes, which are time-uncorrelated random fluctuations occurring independently between multiple observations, and (ii) red noise processes, which characterise a correlated behavior over long timescales, manifesting as slow, systematic variations in the timing residuals. The red noise itself can be further classified as achromatic (time-correlated but observing frequency independent) and chromatic (time-correlated and observing frequency dependent) red noise processes. Amongst the latter one, there is a further sub-division based on the chromatic signature imprinted in the timing residuals – (i) DM noise (varying as $\nu^{-2}$), and (ii) Scattering noise (varying as $\nu^{-4}$ for Gaussian inhomogeneities). In addition to these two dominant noise processes, there may be other chromatic noise processes in pulsars as well.
Apart from these single-pulsar noise processes, one comes across time-correlated signals present in an ensemble of pulsars observed by a PTA. These signals are modeled as common processes for the entire PTA in the form of CRN. This process either does not exhibit any spatial correlations between pulsars and is characterised only by its spectral properties (CURN), or represents the GWB with spatial correlations, depending only on the sky-separations of the pulsar pairs, as described by the Hellings-and-Downs overlap reduction function (Hellings1983) for a GWB stemming from a population of circular inspiraling SMBHBs. Estimation of CURN in PTA datasets informs about the presence of a common red noise process in the ensemble, which may or may not manifest as a GWB Chen2021; Arzoumanian2020; Goncharov2021; Antoniadis2022.

### v2

@@ 1 Introduction
Millisecond pulsars (MSPs) are rapidly spinning neutron stars with extraordinary rotational stability and act as highly accurate cosmic clocks (Lorimer and Kramer, 2004; Hobbs et al., 2019). Pulsar timing is a powerful technique that involves careful monitoring of the precise rotational behavior of MSPs by comparing the measured pulse Times of Arrival (ToAs) with timing models that encapsulate the pulsar’s rotational, astrometric and orbital properties (Edwards et al., 2006). Gravitational Waves (GWs) passing across the line-of-sight to a pulsar modulate the ToAs of the pulsar measured at a terrestrial observatory (Estabrook and Wahlquist, 1975). The ultra-low frequency (nHz) GWs can be produced by populations of inspiraling Supermassive Black Hole Binaries (SMBHBs) (Sesana et al., 2008), cosmic string networks (Kibble, 1976), relic post-inflation GWs (Lasky et al., 2016), cosmological phase transitions (Grishchuk, 2005) or exotic dark matter species (Burke-Spolaor et al., 2019) (and references therein). The incoherent superposition of these nHz GWs in the universe produces a stochastic Gravitational Wave Background (GWB: Mingarelli et al., 2017; Burke-Spolaor et al., 2019; Sesana and Figueroa, 2025). This extremely feeble background introduces tiny delays in the ToAs that depend on the underlying population properties of its GW source. For an isotropic distribution of a population of circular inspiraling SMBHBs, the GWB has a characteristic spectral signature and manifests itself as quadrupolar spatial correlations described by the Hellings-Downs (HD) curve (Hellings and Downs, 1983), and depends only on the relative sky separations of pulsar pairs in the pulsar ensemble used for this purpose.
@@
Due to the high rotational stability of MSPs, such signatures can be detected by precise timing of these pulsars through Pulsar Timing Arrays (PTAs: Sazhin, 1978; Foster and Backer, 1990; Detweiler, 1979; Taylor, 2021; Shaifullah, 2025). PTAs monitor an ensemble of MSPs and leverage the spatially correlated modulations in ToAs to detect and possibly characterize the GWB. However, this signature can be obscured by a variety of noise sources which may be intrinsic to the pulsar, or could be picked up as the pulsed radio emission propagates through the ionized interstellar medium (IISM) before it is received by the terrestrial observatories (Shannon and Cordes, 2010; Verbiest and Shaifullah, 2018; Donner et al., 2020). Careful modeling and characterization of these noise sources, therefore, becomes central to the detection of a GWB. This pursuit of unveiling the existence of a GWB has led to the formation of PTA consortia across the globe, such as the European Pulsar Timing Array (EPTA: Janssen et al. 2008), the Parkes Pulsar Timing Array (PPTA: Manchester et al. 2013), the North American Nanohertz Observatory for Gravitational Waves (NANOGrav: (Jenet et al., 2009; McLaughlin, 2013)), the Indian Pulsar Timing Array (InPTA: Joshi et al. (2018, 2022)), the MeerKAT Pulsar Timing Array (MPTA: Miles et al. 2023), and the Chinese Pulsar Timing Array (CPTA: Lee 2016). Recently, the African Pulsar Timing (APT [^1] https://africanpulsartiming.github.io/) group has also emerged and is engaged in broader pulsar science goals, including PTA science, in collaboration with the MPTA. Collectively, the NANOGrav, InPTA, EPTA, PPTA and APT consortia form the International Pulsar Timing Array (IPTA: Verbiest et al. 2016; Perera et al. 2019) consortium, which combines data and resources of the constituent PTAs to enhance the sensitivity to detect these nHz GWs.
@@
In an earlier study, Chen et al. (2021); Arzoumanian et al. (2020); Goncharov et al. (2021b); Antoniadis et al. (2022) reported the presence of a Common Red Noise (CRN) process in multiple PTA datasets. However, this signal could not be characterized as a GWB due to the absence of statistically significant spatial correlations, even though the process showed temporal correlations, necessary for a common red process. More recently, the NANOGrav, PPTA, CPTA, MPTA and EPTA+InPTA analyses have indicated the presence of a spatially correlated GWB-like signal (Agazie et al., 2023a; Reardon et al., 2023a; Miles et al., 2025b; Xu et al., 2023; Antoniadis et al., 2023), thus ushering the era of ultra-low frequency GW astronomy. However, accurate noise modeling is critical to bring out these spatial correlations and estimate the parameters of the spectral signature of the GWB.
@@
The present study focuses on the customised Single-Pulsar Noise Analysis (SPNA) of 27 MSPs from the InPTA Data Release 2 (Rana et al., 2025b), building up on the work presented in Srivastava et al. (2023), employing the methods of Bayesian inference to detect and quantify the various noise processes underlying PTA datasets. The InPTA dataset has the niche characteristic of simultaneous dual-band observations, which gives it a unique standpoint amongst global PTAs for precise modeling of chromatic processes. The rest of the paper is structured as follows. In Section 2, we give a brief description of the InPTA-DR2 observing strategy, timing analysis and data products. In Section 3, we discuss the noise processes and their modeling in detail, whereas in Section 3.2 we give the theoretical background of noise models and in Section 3.3 we discuss the analysis methodology adopted in the present work. We present the results in Sections 4 and 5, followed by a discussion in Section 6. We conclude with a brief summary in Section 7 and outline the future prospects related to this work in Section 8.
@@ 2 Dataset
The InPTA-DR2 dataset consists of 27 MSPs observed with a cadence of $\sim$$15$ days using the upgraded Giant Metrewave Radio Telescope (uGMRT: Swarup, 1991; Gupta et al., 2017), which enables simultaneous multi-band observations. The dataset spans $\sim$$7.2$ years of total time baseline ($T_{\text{span}}$) – nearly twice that of InPTA DR1 (Tarafdar et al., 2022), and includes concurrent observations in the $300-500\,\text{MHz}$ (Band 3) and $1260-1460\,\text{MHz}$ (Band 5) frequency ranges, with $100\,\text{MHz}$ or $200\,\text{MHz}$ bandwidths in each band, depending on the observing epoch. This leads to a maximum of seven different backend-receiver configurations for the longest time baseline pulsars. The raw data was recorded with the GMRT Wideband Backend and observations employed the phased-array (PA) beam and the real-time coherent de-dispersion pipeline (CDP) in one or both bands, depending on the observing cycle (see Table 1 of Rana et al., 2025a). The ‘Pipeline for the Indian Pulsar Timing Array’ (PINTA: Susobhanan et al., 2021) is used to perform RFI mitigation and partial folding, to convert the uGMRT raw data into a standard pulsar archive format (PSRFITS: Hotan et al., 2004) with full frequency resolution and several sub-integrations of $10\,\text{s}$ duration. Further technical details of InPTA observations and data reduction are provided in Sections 2 & 3 and illustrated in Figures 1 & 2 of Rana et al. (2025a).
@@
To generate noise-free templates in each frequency band – (i) high signal-to-noise ratio (S/N) observations (free of artifacts) are selected in both bands, (ii) a preliminary alignment of the pulse profiles is performed using an initial fiducial dispersion measure [^2] The integrated free electron column density along the line-of-sight to a pulsar is called the Dispersion Measure (DM), given by : $\text{DM}=\int_{\ell}n_{e}\,\mathrm{d}l$ (DM), (iii) the bandpass shape is corrected by equalizing the off-pulse RMS across frequency channels, and (iv) optimal wavelet smoothing is applied using the psrsmooth tool of PSRCHIVE (Hotan et al., 2004). The final templates are then de-dispersed using a high-precision fiducial DM obtained through an iterative procedure. This approach minimizes pulse-profile smearing in template profiles due to uncorrected DM delays. Furthermore, a refined procedure is employed to determine the optimal number of frequency sub-bands to partially collapse the data across frequency for each pulsar in the two bands. This optimization balances two requirements – (i) to provide sufficient S/N and similar ToA precision in each sub-band, and (ii) to minimize radio-frequency dependent profile evolution. Details about these aspects are provided in Sections 4.1–4.3 of Rana et al. (2025a).
@@
Given the noise-free templates, optimized number of sub-bands, and fiducial DM estimates for each pulsar, the DMCalc (Krishnakumar et al., 2021) pipeline is used to estimate frequency-resolved ToAs and DMs, along with DMXs [^3] DMXs are defined as the differences between the estimated DMs and fiducial DM, and correspond to a linear piece-wise fit to the DM model in time-domain, to account for temporal variations in the DM. computed from the estimated and fiducial DMs for every epoch. In the deterministic timing analysis, these epoch-wise DMXs are used to model dispersion delays. ToAs with large uncertainties or having poor S/N in a given sub-band and/or affected by radio-frequency interference (RFI) are flagged. Finally, the timing model parameters are fitted using the TEMPO2 package (Hobbs et al., 2006). Following these procedures, the resulting InPTA-DR2 dataset spans MJD 577686039957768-60399 (uGMRT observation cycles 314531-45, excluding cycle $36$), where the exact temporal coverage differs for each pulsar. However, the dataset for PSR J1713++0747 contains only the pre-event ToAs (Singha et al., 2021) leading to a reduced time baseline for this pulsar. Further details of the DMCalc pipeline and timing analysis are provided in Sections 4.4 and 5 of Rana et al. (2025a), respectively. The final InPTA-DR2 data products included Band $3+5$ and Band $3$ DM time series, par files containing the deterministic timing solutions with DMXs, and tim files containing the cleaned ToAs [^4] https://github.com/inpta/InPTA.DR2.git.
@@ 3 Noise modeling methodology
ToAs obtained from the pulsed emission of pulsars depend on a variety of deterministic and stochastic processes. The deterministic contributions are derived from the rotational, astrometric and orbital dynamics of the pulsar, the orbital motion of the Earth, as well as the propagation effects of the interstellar medium. These deterministic terms are captured by the timing model of the pulsar derived from the pulsar timing ephemeris. However, the stochastic components cannot be accurately accounted for, based on this deterministic timing model alone. These components include achromatic (such as spin noise) as well as chromatic (such as DM noise) processes that add to the overall measurement noise in the dataset. Inaccurate modeling of any of these noise processes may lead to false detections and incorrect characterization of potential GW signals. Therefore, precise modeling of these processes is of immense importance in PTA data analyses to isolate the astrophysical signals of interest (Di Marco et al., 2025). In this section, we describe, in detail, the theoretical construct of various deterministic and stochastic noise processes and the customized SPNA methodology employed to estimate their contributions to the overall InPTA-DR2 noise budget.
@@ 3.1 The Deterministic Timing Model
The deterministic timing model essentially provides a mapping of a ToA measured at the terrestrial observatory to a fiducial point on the emitted pulse from a pulsar by performing a series of transformations from the pulsar frame to an inertial frame at the Solar System Barycenter (SSB). These transformations include clock corrections to convert the local observatory time to the Universal Time ($\Delta_{\text{clock}}$: Panfilo and Arias, 2019), account for the arrival time delays due to differences in the topocentric and SSB reference frames ($\Delta_{\odot}$: Park et al., 2021), the constant light-travel time between the pulsar and the Earth ($\Delta_{\gamma}$), propagation through the IISM ($\Delta_{\text{IISM}}$), the astrometric and binary orbital motion of the pulsar ($\Delta_{\text{binary}}$). Therefore, the observed arrival time ($t_{\text{obs}}$) of the pulse can be related to the emission time ($t_{\text{em}}$) as:
@@
These delays are incorporated in the deterministic timing model in the form of a par file which is then used to perform standard timing analysis using packages such as TEMPO2 (Hobbs et al., 2006) and PINT (Luo et al., 2021) to generate a best-fit timing solution. This solution is then used to derive the ToA (or timing) residuals, which are essentially the differences in the observed ($t_{\text{obs}}$) ToAs and the predicted ToAs as per the best-fit timing solution ($\bar{t}_{\text{obs}}$). The additive terms $\mathcal{N}$ in equation 1 correspond to the residual noise in the measured ToAs, which can arise due to noise in the radiometer or intrinsic fluctuation of the pulsed emission ($\mathcal{N}_{\text{r}}$) or other stochastic noise processes ($\mathcal{N}_{\text{s}}$), such as the spin noise or DM noise, etc., that are not modeled completely by the deterministic timing model of the pulsar.
@@ 3.2 The Noise Processes
Based on the temporal properties of noise, the single-pulsar noise processes are categorized as (i) white noise processes, which are time-uncorrelated random fluctuations occurring independently between multiple observations, and (ii) red noise processes, which characterize a correlated behavior over long timescales, manifesting as slow, systematic variations in the timing residuals. The red noise itself can be further classified as achromatic (time-correlated but observing frequency independent) and chromatic (time-correlated and observing frequency dependent) red noise processes. Amongst the latter one, there is a further sub-division based on the chromatic signature imprinted in the timing residuals – (i) DM noise (varying as $\nu^{-2}$), and (ii) Scattering noise (varying as $\nu^{-4}$ for Gaussian inhomogeneities). In addition to these two dominant noise processes, there may be other chromatic noise processes in pulsars as well.
Apart from these single-pulsar noise processes, one encounters time-correlated signals present across the entire ensemble of pulsars observed by a PTA. These signals are modeled as common processes for the entire PTA in the form of CRN. This process either does not exhibit any spatial correlations between pulsars and is characterized only by its spectral properties (referred to as the Common Uncorrelated Red Noise or CURN), or represents the spatial correlations between pulsar pairs that depend only on their angular separations on the sky. These spatial correlations can exist in a variety of modes, such as monopole, dipole, quadrupole, etc., each of which could have multiple astrophysical origins (Sesana and Figueroa, 2025). One particular form of quadrupolar spatial correlations are described by the Hellings-and-Downs overlap reduction function (Hellings and Downs, 1983) for a GWB stemming from a population of circular inspiraling SMBHBs. The presence of such quadrupolar correlations are of special interest in PTA science and are considered as the smoking gun for GWB detection (Agazie et al., 2023a; Reardon et al., 2023a; Miles et al., 2025b; Xu et al., 2023; Antoniadis et al., 2023). In reality, one can expect a superposition of CURN and GWB signals, therefore, the estimation of CURN in PTA datasets informs about the presence of a common red noise process in the ensemble, which may or may not manifest as a GWB Chen et al. (2021); Arzoumanian et al. (2020); Goncharov et al. (2021b); Antoniadis et al. (2022).